OpenStudy (anonymous):

@agent0smith How would you factor an expression when the expression starts with a number instead of a variable?

OpenStudy (agent0smith):

I thought you were posting an actual question... with numbers and all

OpenStudy (anonymous):

The expression im trying to solve is 10y^2+39y+14.

OpenStudy (agent0smith):

10y^2+39y+14 First step is to multiply the first and last numbers: 10*14 = 140 now find two numbers that multiply to 140 and add up to 39

OpenStudy (anonymous):

I got the first part. Kinda confused on the second.

OpenStudy (agent0smith):

find two numbers that multiply to 140 and add up to 39?

OpenStudy (anonymous):

Yes.

OpenStudy (agent0smith):

I'm not sure what confuses you...

OpenStudy (anonymous):

I'm sorry. I'm just really bad at math and i don't understand what you mean when you say that.

OpenStudy (anonymous):

You will have to explain it more.

OpenStudy (agent0smith):

This is the same method you used yesterday...

OpenStudy (agent0smith):

find two numbers that when you multiply them, it equals 140, and when you add them, it equals 39

OpenStudy (anonymous):

oh. ok. i thought when you said that you meant something that just didnt make any sense.

OpenStudy (agent0smith):

140 = 35*4 and 35 + 4 = 39 so those are your two numbers. Now the next step...